# The Algebra of the Pseudo-Observables I: Why Quantum Mechanics is the ultimate description of Reality

### v5

@@ 1.2 The quest for a new formulation
In the meantime, in 1952, David Bohm, dissatisfied with the prevailing orthodoxy, developed a theory[32, 33] that is essentially the same as de Broglie’s pilot wave theory, the only difference being that de Broglie’s dynamics is formulated in terms of velocity rather than acceleration. However the general interest in the theory was little, till when, in 1966, John Bell[34] rediscovered the results of Grete Hermann’s work, i.e. that “von Neumann’s no-hidden-variables proof was based on an assumption that can only be described as silly”[35]. The von Neumann argumentation, in facts, demonstrates only that hidden-variables theories must be nonlocal.
@@
Another interesting line of research in the field of quantum mechanics formulations is that of algebraic theories[40]. Algebraic quantum mechanics is an abstraction and generalization of the Hilbert space formulation. In the development of the algebraic approach a major role was played by von Neumann. His joint paper with Jordan and Wigner[41] was, in facts, one of the first attempts to go beyond Hilbert space of state vectors. Subsequently, von Neumann developed the mathematical theory of operator algebras in a series of outstanding papers[42, 43, 44, 45, 46]. Notwithstanding his efforts, von Neumann’s attempts to apply this theory to quantum mechanics were unsuccessfully[47]. The operator algebras that he introduced, now called von Neumann algebras, however, still play a central role in the algebraic approach to quantum theory.
@@ 1.3 What it has been done in this paper
Proceeding in such a manner, every aspect of quantum mechanics acquires a clear physical interpretation or a logical explanation, providing, for instance, in a natural way the reason for the structure of complex algebra and the matrix structure of Werner Heisenberg’s formulation of quantum mechanics.
@@ 4.1 Dyadic forms
$C=1\,C\,1=\sum_{j,k}I_{j}CI_{k}\text{ .}$ (35)
According to (35), any observable $C$ is then given by a linear combination of the pseudo-observables obtained by multiplying the observable to the right and to the left for an element of the basis of pairwise orthogonal primitive projectors. This result can be immediately extended to pseudo-observables also, by virtue of (22). It can thus be said that the space of the pseudo-observables is spanned by the set $\left\{I_{j}CI_{k}\right\}$, where one should consider the products for every observable $C$. However, in fact, most of the elements of this set are linearly dependent on the others.
Initially we consider a product of the form: $I_{j}CI_{j}$, that will be called, for a reason that will be explained shortly, projection of $C$ according the projector $I_{j}$. It is immediate, first of all, to prove that this expression is an observable that commutes with each of the primitive projectors of the basis $\left\{I_{k}\right\}$. By the theorem proved at the end of subsection 3.4, the projection is expressible as a linear combination of the projectors of the basis $\left\{I_{k}\right\}$. Applying the associative property of the product of observables and the fact that the projectors of the basis are pairwise orthogonal, it is concluded finally that the projection of $C$ according the projector $I_{j}$ is proportional to the only projector $I_{j}$:
@@
It is interesting to observe that the relation (36) is characteristic of primitive projectors. Be, in fact, $I$ a projector such that for each observable $C$ it results:
@@
Let us now consider a product of the form: $I_{j}CI_{k}$, where $C$ is a generic observable. It will be called dyadic form relative to the pair of primitive projectors $\left(I_{j},I_{k}\right)$ having core $C$. We state the following postulate on the dyadic forms [^4] In the Dirac formulation of quantum mechanics, the dyadic forms correspond to operators of the type: $\gamma\,|j\rangle\langle k|$, where the indexes $j$ and $k$ denote distinct eigenvectors, normalized to one, of a certain linear Hermitian operator and $\gamma$ is a suitable complex constant, dependent on the core. If $\hat{C}$ is the operator associated to the observable $C$, in fact, it results: $\gamma=\langle j|\hat{C}|k\rangle$.: considered a pair of primitive projectors $\left(I_{j},I_{k}\right)$ there is always an observable $C_{jk}$ such that the dyadic form relative to the pair of primitive projectors and having core $C_{jk}$ is not the zero observable; furthermore, given a whatever observable $C$, there exists a complex constant $\alpha$ such that it is:
$I_{j}CI_{k}=\alpha\,I_{j}C_{jk}I_{k}\text{ .}$ (39)
It is worth observing that, in the case where the indexes $j$ and $k$ are equal, the postulate on the dyadic forms reduces itself to what before has been proved about the projections according primitive projectors. Note, also, that the relation (39) is trivially satisfied if the observable $C$ is compatible, and therefore commutes, with $I_{j}$ or with $I_{k}$, because, if $j\neq k$, it results $I_{j}CI_{k}=0$. The postulate, therefore, actually concerns only those cases in which we consider products between incompatible observables that do not have as result an observable, for which, up to this point, there was provided no actual definition.
The postulate on the dyadic forms, therefore, completes the definition of the product of pseudo-observables, compared to what was done in subsection 2.4, for compatible observables. This definition allows to minimize the number of not actually measurable pseudo-observables in the space $\mathbb{P}$ and therefore represents a further application of the Occam’s razor principle.
By virtue of the postulate on the dyadic forms and according to (35) we can now state that the pseudo-observable space $\mathbb{P}$ is spanned by the set of dyadic forms $\left\{\Phi_{jk}:=I_{j}C_{jk}I_{k}\right\}$, where the observables $C_{jk}$ are chosen in such a way that they make not null the dyadic forms of which they are core. If, in fact, $P$ is a generic pseudo-observable, whose real part is $P_{\mathrm{R}}$ and the imaginary part is $P_{\mathrm{I}}$, one has:
$P=\sum_{j,k}\left(I_{j}P_{\mathrm{R}}I_{k}+iI_{j}P_{\mathrm{I}}I_{k}\right)\text{ .}$ (40)
For each pair of indexes $j$ and $k$, according to (39), exist two complex constants, $\alpha_{jk}$ and $\beta_{jk}$, such that it results:
$I_{j}P_{\mathrm{R}}I_{k}=\alpha_{jk}\,\Phi_{jk}\qquad\mbox{and}\qquad I_{j}P_{\mathrm{I}}I_{k}=\beta_{jk}\,\Phi_{jk}$
which, once replaced in (40), allow to obtain:
$P=\sum_{j,k}\left(\alpha_{jk}+i\beta_{jk}\right)\Phi_{jk}=\sum_{j,k}\varpi_{jk}\,\Phi_{jk}$ (41)
where we have put: $\varpi_{jk}:=\alpha_{jk}+i\beta_{jk}$.
The dyadic forms of the set $\left\{\Phi_{jk}\right\}$ are, also, linearly independent. If, in fact, we consider a linear combination of them, by means of a set of complex constants $\left\{\alpha_{jk}\right\}$, and set it to be equal to the zero observable:
$\sum_{j,k}\alpha_{jk}\,\Phi_{jk}=0$ (42)
$\alpha_{j^{\prime}k^{\prime}}\,\Phi_{j^{\prime}k^{\prime}}=0\;\Rightarrow\;\alpha_{j^{\prime}k^{\prime}}=0$
from which follows the linear independence of the dyadic forms.
The set of dyadic forms $\left\{\Phi_{jk}\right\}$ therefore constitutes a basis, that, in our assumptions, is finite or countable to the limit, for the pseudo-observable space. This basis is equipotent to the Cartesian product of the basis of primitive projectors by itself. Since the bases of a vector space are mutually equipotent, an important consequence of this result is that also the bases of primitive projectors associated with the various spaces of compatible observables must be each other equipotent.
Since the set of dyadic forms $\left\{\Phi_{jk}\right\}$ constitutes a basis for the space $\mathbb{P}$, each coefficient $\varpi_{jk}$ of the decomposition (41) of a generic pseudo-observable $P$ is, therefore, uniquely determined and will be called component of the pseudo-observable according the dyadic form $\Phi_{jk}$.
It is now worth observing that if $K$ is a generic pseudo-observable such that the product $I_{j}KI_{k}$ is not null, according to (41), it must exist a non-zero complex constant $\kappa_{jk}$ such that it results $I_{j}KI_{k}=\kappa_{jk}\,\Phi_{jk}$ and therefore:
$\Phi_{jk}=I_{j}\left(\frac{1}{\kappa_{jk}}K\right)I_{k}$
that implies that one can safely assume that the cores of the dyadic form are pseudo-observables.
Note, finally, that, if $k\neq l$, by virtue of (6), it is always:
$\Phi_{jk}\Phi_{lm}=0\text{ .}$ (43)
@@ 4.2 Dyad bases
We will now prove that it is possible to choose the dyadic form pseudo-observable cores of the basis in such a manner to obtain a set of normalized dyadic forms or dyads [^5] Dyads correspond in the Dirac formulation of quantum mechanics to operators of the form: $|j\rangle\langle k|$, where the indexes $j$ and $k$ denote distinct eigenvectors normalized to one, of a given linear Hermitian operator. The name of dyad will find justification in subsection “State vectors” of the second paper., that will be denoted by the symbol $\Gamma_{jk}$, satisfying the following conditions:
@@
and the property immediately follows on the basis of the orthogonality relation (43).
The dyad bases are the natural extensions to the whole space of pseudo-observables of the bases of pairwise orthogonal primitive projectors and allow one to make explicit, in terms of components, the operations of sum and of product of pseudo-observables, completing, as already stated, their definition. Let, in fact, be $P$ a pseudo-observable and $\left\{\Gamma_{jk}\right\}$ a dyad basis. By decomposing the pseudo-observable $P$ according the dyad basis, as in (41), one has:
$P=\sum_{j,k}\varpi_{jk}\,\Gamma_{jk}$ (49)
where the coefficients $\varpi_{jk}$, the dyadic components, are suitable complex constants. The components $\varpi_{jk}$ may be thought, in the finite-dimensional cases, as elements of a matrix that will be called matrix associated to the pseudo-observable with respect to the dyad basis. On the basis of this interpretation, we will call diagonal components the dyadic components of the form $\varpi_{jj}$.
@@
By multiplying both sides of (49) for the primitive projectors of the basis from which the dyads are deduced, one obtains the following important relation:
$I_{j}PI_{k}=\varpi_{jk}\,\Gamma_{jk}$ (51)
that generalizes relation (10) and supplies a characterization of the components of a pseudo-observables with respect to a dyad basis.
@@
$Q=\sum_{j,k}\vartheta_{jk}\,\Gamma_{jk}$
where $\vartheta_{jk}$ are suitable complex constants, one has:
$P+Q=\sum_{j,k}\left(\varpi_{jk}+\vartheta_{jk}\right)\Gamma_{jk}$ (52)
@@
$PQ=\sum_{j,k}\left(\sum_{l}\varpi_{jl}\,\vartheta_{lk}\right)\Gamma_{jk}$ (53)
@@
It will be, now, clarified better the tie between dyad basis and basis of pairwise orthogonal primitive projectors. We will say, first of all, that a dyad basis $\left\{\Gamma_{jk}\right\}$ is associated to the basis $\left\{I_{j}\right\}$ of pairwise orthogonal primitive projectors, if all the dyadic forms of the first are relative to pairs of projectors of the basis $\left\{I_{j}\right\}$. This stated, let $\left\{\Gamma_{jk}\right\}$ and $\left\{\tilde{\Gamma}_{jk}\right\}$ be two dyad bases associated to the same basis $\left\{I_{j}\right\}$ of pairwise orthogonal primitive projectors. By virtue of the postulate on the dyadic forms, introduced in subsection 4.1, there must exist complex constants $\gamma_{jk}$ such that it results:
[…]
The second property of dyads implies:
@@
By virtue of this relation and of the second and third property of the dyads, one, besides, has:
$I_{j}=\tilde{\Gamma}_{jk}\left(\tilde{\Gamma}_{jk}\right)^{\dagger}=\left|\gamma_{jk}\right|^{2}I_{j}\;\Rightarrow\;\left|\gamma_{jk}\right|^{2}=1$
@@
$\tilde{\Gamma}_{jk}=\tilde{\Gamma}_{jk_{0}}\left(\tilde{\Gamma}_{kk_{0}}\right)^{\dagger}=e^{i\left(\vartheta_{j}-\vartheta_{k}\right)}\Gamma_{jk}$ (59)
@@ 5.1 Summary of Postulates
    It is, last, assumed the validity of postulate on the dyadic forms, according to which, for each pair of primitive projectors of a basis, there is a non-zero associated dyadic form; every other dyadic form associated with the same pair of primitive projectors is obtained by multiplying that dyadic form by a suitable complex constant (i.e., dyadic forms relative to the same pair of primitive projectors of the same basis are linearly dependent on each other). The postulate on the dyadic forms completes, in a coherent manner, the definition of product among non compatible observables and allows to extend the bases of pairwise orthogonal primitive projectors, that span complete subspaces of compatible observables, to dyad bases, that span the whole space $\mathbb{P}$ of the pseudo-observables. The decomposition of the pseudo-observables according the dyads of a basis allows, finally, one to write explicit expressions for sums and products of observables, compatible or not, or pseudo-observables.

### v6

@@ 1.2 The quest for a new formulation
In the meantime, in 1952, David Bohm, dissatisfied with the prevailing orthodoxy, developed a theory[32, 33] that is essentially the same as de Broglie’s pilot wave theory, the only difference being that de Broglie’s dynamics is formulated in terms of velocity rather than acceleration. However the general interest in the theory was little, till when, in 1966, John Bell[34] rediscovered the results of Grete Hermann’s work, i.e. that “von Neumann’s no-hidden-variables proof was based on an assumption that can only be described as silly”[35]. The von Neumann argumentation, in fact, demonstrates only that hidden-variables theories must be nonlocal.
@@
Another interesting line of research in the field of quantum mechanics formulations is that of algebraic theories[40]. Algebraic quantum mechanics is an abstraction and generalization of the Hilbert space formulation. In the development of the algebraic approach a major role was played by von Neumann. His joint paper with Jordan and Wigner[41] was, in fact, one of the first attempts to go beyond Hilbert space of state vectors. Subsequently, von Neumann developed the mathematical theory of operator algebras in a series of outstanding papers[42, 43, 44, 45, 46]. Notwithstanding his efforts, von Neumann’s attempts to apply this theory to quantum mechanics were unsuccessfully[47]. The operator algebras that he introduced, now called von Neumann algebras, however, still play a central role in the algebraic approach to quantum theory.
@@ 4.1 Projections and Dyadic Forms
$P=1\,P\,1=\sum_{j,k}I_{j}PI_{k}\text{ .}$ (35)
According to (35), any pseudo-observable $P$ is then given by a linear combination of the pseudo-observables obtained by multiplying the pseudo-observable to the right and to the left for an element of the basis of pairwise orthogonal primitive projectors.
We now consider a product of the form: $I_{j}CI_{j}$, where $C$ is an observable, that will be called, for a reason that will be explained shortly, projection of $C$ according the projector $I_{j}$. It is immediate, first of all, to prove that this expression is an observable that commutes with each of the primitive projectors of the basis $\left\{I_{k}\right\}$. By the theorem proved at the end of subsection 3.4, the projection is expressible as a linear combination of the projectors of the basis $\left\{I_{k}\right\}$. Applying the associative property of the product of observables and the fact that the projectors of the basis are pairwise orthogonal, it is concluded finally that the projection of $C$ according the projector $I_{j}$ is proportional to the only projector $I_{j}$:
@@
It is interesting to observe that the relation (36) is characteristic of primitive projectors. Let, in fact, $I$ be a projector such that for each observable $C$ it results:
@@
Projections are easily extended also to the case of pseudo-observables. In fact, if $P\in\mathbb{P}$ is a pseudo-observable and $I$ is a primitive projector, by making use of (22), one immediately obtains:
$IPI=IP_{\mathrm{R}}I+i\,IP_{\mathrm{I}}I=(p_{\mathrm{R}}+ip_{\mathrm{I}})I=\varpi I$ (39)
where $p_{\mathrm{R}}$ and $p_{\mathrm{I}}$ are suitable real constants and $\varpi:=p_{\mathrm{R}}+i\,p_{\mathrm{I}}$ a complex constant.
We now define dyadic form relative to the pair of primitive projectors $\left(I_{j},I_{k}\right)$ with core $P$ the product: $I_{j}PI_{k}$, where $P$ is a generic pseudo-observable. In next subsection, we will be able to construct a useful basis for the space $\mathbb{P}$ of the pseudo-observables formed by dyadic forms with suitably chosen cores.
To this aim, it will be assumed that for each pair of primitive projectors $\left(I_{j},I_{k}\right)$ there is always a pseudo-observable $P_{jk}$ such that the dyadic form relative to the pair of primitive projectors with core $P_{jk}$ is nonzero.
@@ 4.2 Dyad bases
We will now prove that it is possible to construct a set of dyadic form with suitably chosen pseudo-observable cores such that it constitutes a basis for the whole space $\mathbb{P}$of the pseudo-observables. Such dyadic forms will be called dyads [^4] Dyads correspond in the Dirac formulation of quantum mechanics to operators of the form: $|j\rangle\langle k|$, where the indexes $j$ and $k$ denote distinct eigenvectors normalized to one, of a given linear Hermitian operator. The name of dyad will find justification in subsection “State vectors” of the second paper., here denoted by the symbols $\Gamma_{jk}$, and satisfy the following conditions:
@@
and the property immediately follows if one consider also the orthogonality relation between the projectors of the basis.
We can now prove an important property of the dyadic forms. For each pseudo-observable $P$, let’s consider the dyadic form relative to the pair of primitive projectors $\left(I_{j},I_{k}\right)$ with core $P$. By making use of (39) and of the properties of the dyads, one has:
$I_{j}PI_{k}=I_{j}P\,\Gamma_{kj}\Gamma_{jk}=\left(I_{j}\left(P\,\Gamma_{kj}\right)I_{j}\right)\Gamma_{jk}=(\varpi_{jk}\,I_{j})\Gamma_{jk}=\varpi_{jk}\Gamma_{jk}$ (45)
that generalizes the projections.
By virtue of the postulate on the dyadic forms and we can now state that the pseudo-observable space $\mathbb{P}$ is spanned by the set of dyads $\left\{\Gamma_{jk}\right\}$. If, in fact, $P$ is a generic pseudo-observable, by making use of the decomposition (35) and of (45), it results:
$P=\sum_{j,k}I_{j}PI_{k}=\sum_{j,k}\varpi_{jk}\,\Gamma_{jk}$ (46)
where the coefficients $\varpi_{jk}$, the dyadic components, are suitable complex constants.
Dyads are, also, linearly independent. If, in fact, we consider a linear combination of them, by means of a set of complex constants $\left\{\alpha_{jk}\right\}$, and set it to be equal to the zero observable:
$\sum_{j,k}\alpha_{jk}\,\Gamma_{jk}=0$ (47)
$\alpha_{j^{\prime}k^{\prime}}\,\Gamma_{j^{\prime}k^{\prime}}=0\;\Rightarrow\;\alpha_{j^{\prime}k^{\prime}}=0$
from which follows the linear independence of the dyads.
The set of dyads $\left\{\Gamma_{jk}\right\}$ therefore constitutes a basis, that, in our assumptions, is finite or countable to the limit, for the pseudo-observable space. This basis is equipotent to the Cartesian product of the basis of primitive projectors by itself. Since the bases of a vector space are mutually equipotent, an important consequence of this result is that also the bases of primitive projectors associated with the various spaces of compatible observables must be each other equipotent.
Since the set of dyads $\left\{\Gamma_{jk}\right\}$ constitutes a basis for the space $\mathbb{P}$, each coefficient $\varpi_{jk}$ of the decomposition(46) of a generic pseudo-observable $P$ is, therefore, uniquely determined and will be called component of the pseudo-observable according to the dyad $\Gamma_{jk}$.
The dyad bases are therefore t natural extensions to the whole space of pseudo-observables of the bases of pairwise orthogonal primitive projectors and allow one to make explicit, in terms of components, the operations of sum and of product of pseudo-observables, completing, as already stated, their definition.
The components $\varpi_{jk}$ may be thought, in the finite-dimensional cases, as elements of a matrix that will be called matrix associated to the pseudo-observable with respect to the dyad basis. On the basis of this interpretation, we will call diagonal components the dyadic components of the form $\varpi_{jj}$.
@@
$Q=\sum_{j,k}\theta_{jk}\,\Gamma_{jk}$
where $\theta_{jk}$ are suitable complex constants, one has:
$P+Q=\sum_{j,k}\left(\varpi_{jk}+\theta_{jk}\right)\Gamma_{jk}$ (49)
@@
$PQ=\sum_{j,k}\left(\sum_{l}\varpi_{jl}\,\theta_{lk}\right)\Gamma_{jk}$ (50)
@@
It will be, now, clarified better the tie between dyad basis and basis of pairwise orthogonal primitive projectors. We will say, first of all, that a dyad basis $\left\{\Gamma_{jk}\right\}$ is associated to the basis $\left\{I_{j}\right\}$ of pairwise orthogonal primitive projectors, if all the dyadic forms of the first are relative to pairs of projectors of the basis $\left\{I_{j}\right\}$. This stated, let $\left\{\Gamma_{jk}\right\}$ and $\left\{\Gamma^{\prime}_{jk}\right\}$ be two dyad bases associated to the same basis $\left\{I_{j}\right\}$ of pairwise orthogonal primitive projectors. By virtue of the (46), there must exist complex constants $\gamma_{jk}$ such that it results:
$\Gamma^{\prime}_{jk}=\gamma_{jk}\,\Gamma_{jk}\text{ .}$ (51)
The second property of dyads, besides, implies:
@@
By virtue of this relation and of the second and third property of the dyads, one, therefore, has:
$I_{j}=\Gamma^{\prime}_{jk}\left(\Gamma^{\prime}_{jk}\right)^{\dagger}=\left|\gamma_{jk}\right|^{2}I_{j}\;\Rightarrow\;\left|\gamma_{jk}\right|^{2}=1$
@@
$\Gamma^{\prime}_{jk}=\Gamma^{\prime}_{jk_{0}}\left(\Gamma^{\prime}_{kk_{0}}\right)^{\dagger}=e^{i\left(\vartheta_{j}-\vartheta_{k}\right)}\Gamma_{jk}$ (56)
@@ 5.1 Summary of Postulates
    It is, last, assumed that, for each pair of primitive projectors of a basis, there is a nonzero associated dyadic form. This postulate allows to extend the bases of pairwise orthogonal primitive projectors, that span complete subspaces of compatible observables, to dyad bases, that span the whole space $\mathbb{P}$ of the pseudo-observables. The decomposition of the pseudo-observables according the dyads of a basis allows, finally, one to write explicit expressions for sums and products of observables, compatible or not, or pseudo-observables.